OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 12x - y = -4 4x - 3y = -6

OpenStudy (anonymous):

A= 4x - 3(-12x - 4) = -6 B= 4x - 3(12x + 4) = -6 C= 4(12x + 4) - 3y = -6 D= 4(-12x - 4) - 3y = -6

OpenStudy (anonymous):

plz helo me

OpenStudy (anonymous):

look at your first equation 12x - y = -4. Can you solve for y ?

OpenStudy (anonymous):

I really dont even know where to start

OpenStudy (anonymous):

12x - y = -4 --- add y to both sides 12x - y + y = -4 + y --- combine like terms 12x = -4 + y --- now add 4 to both sides 12x + 4 = -4 + 4 + y --- combine like terms 12x + 4 = y are you still with me ?

OpenStudy (anonymous):

now that we know y = 12x + 4, we can sub that in for y in the other equation 4x - 3y = -6 4x - 3(12x + 4) = -6 Understand ?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem :)